# FiniteTimeBibliography.bib

### v1

@@ 8 Discussion: A sketch of experimental measurement protocols
If the $t_{j}$ are sampled irregularly with average spacing $\delta t$ according to some sufficiently random statistical (e.g. Poisson) process, then $\widetilde{v}(E)$ instead becomes quasiperiodic, with a significantly larger “recurrence” energy $E_{\text{rec}}$ (this is connected to quantum recurrence times [QuantumRecurrences, BrownSusskind2] for an irregular energy spectrum, except in our case time and energy have switched places). In particular, if one has $\tau$ such irregular samples of time and $\widetilde{v}(E_{0})=1$, one expects $\widetilde{v}(E)\ll 1$ between $E_{c}$ and $E_{c}\pm E_{\text{rec}}$, where
@@
This means that the number of “extraneous” energy bands or shells is reduced exponentially with the number of samples $\tau$ if one adopts an irregular sampling of times. Often, the “full width” of the energy spectrum in a many-body system (e.g., the separation between the lowest and highest energy levels) scales [ShenkerThouless, BukovSelsPolkovnikov] as a small power of $N$:
@@
A remaining question is whether we can still generate completely positive functions $v_{+}(t)$ of this type. This is indeed possible [^20] For a related argument in a separate context that involves convolutions and applies to real-valued functions, due to Laura Shou, see Ref. [dynamicalqfastscrambling, Appendix D 1]. We use a slightly different argument here to avoid the requirement of real-valuedness, i.e., we do not require $\sum_{k=1}^{\eta}e^{-i\zeta_{k}(E-E_{c})}\in\mathbb{R}$, which allows a more unconstrained choice of the $\zeta_{k}$., and one example is the spectral form factor $K(t)=\lvert L_{\mathcal{H}}(t)\rvert^{2}$ of some energy spectrum, which remains non-negative $K(t)>0$, and whose Fourier transform is the probability distribution (and therefore, non-negative) of two-level spacings in the spectrum. In our case, where the samples are of time rather than energy levels, we should identify an analogue of the spectral form factor with the energy domain and the $2$-level distribution with the time domain. This suggests that, given a set of $\eta$ points $\zeta_{1},\ldots,\zeta_{\eta}$, we can set $\tau=\eta^{2}$ and
@@ 9 Conclusion
We have described an approach to quantum statistical mechanics that is based on the accessible dynamical properties of a single initial state corresponding to an observable instead of the detailed properties of the energy levels, mirroring classical approaches [KhinchinStatMech, GallavottiErgodic] that bypass the ergodic hypothesis. Crucially, as described in Sec. 8, this allows us to make a conclusive experimental determination of thermalization even in large systems of several qubits with tractable resources. While our results center around energy-band thermalization as a finite-resolution variant of eigenstate thermalization, which is aesthetically desirable given the viewpoint that the energy eigenvalues and eigenstates completely determine quantum dynamics, we found that it is possible, and perhaps even more convenient, to directly connect time-domain quantities to thermalization and bypass the energy domain [^21] We view this as being analogous to the “energy-time uncertainty principle” for many-body systems in Refs. [dynamicalqspeedlimit, dynamicalqfastscrambling], where instead of using the spectral form factor to determine the structure of the energy spectrum and then formulate a speed limit in terms of energy parameters, we directly formulate a time-domain speed limit in terms of the time-domain spectral form factor, which implicitly contains information about the energy domain but is best measured in the time domain..
From an analytical standpoint, our results generalize the observed connection between classical ergodicity and eigenstate thermalization in quantized classical systems [Shnirelman, CdV, ZelditchOG, ZelditchTransition, Sunada, ZelditchMixing, Zelditch, Anantharaman] to observable-dependent quantum thermalization in fully quantum systems. In doing so, we have introduced a framework to account for thermalization over finite energy and time scales, in particular showing that these finite-time features are sufficient to determine thermalization over arbitrarily long timescales due to interference effects (which, to our knowledge, has no obvious classical counterpart). In addition to thermalization, we used this approach to obtain a finite-time Mazur-Suzuki inequality for autocorrelators in terms of approximately conserved charges in App. A, which appears to be the first rigorous result of this nature and may be useful for quantum transport problems in finite but large systems. We have also developed intrinsically quantum methods to access thermalization in energy shells and conserved charges that again rely on interference effects with no classical counterpart.
@@
Finally, we address an important question that has been of some concern in the literature, which is the problem of the physical relevance of “weak” thermalization. In particular, our results show that thermalization can be established in an accessible manner for “almost all” (rather than precisely all) physical states (i.e., “weak” thermalization), such as computational basis states. While our methods can also show thermalization in every conceivable initial state (i.e., “strong” thermalization), as discussed in Sec. 4, this requires a level of accuracy that we do not expect to be experimentally accessible (nor analytically tractable except perhaps in special cases [KawamotoStrongETH]) in the thermodynamic limit of many particles. Given that our focus has been on “accessible” aspects of quantum statistical mechanics, it is worth highlighting some common objections to weak thermalization and their relevance to our approach.
@@
Let us then consider “weak ETH”, which refers to the diagonal part of ETH [Eq. (3)] applying to almost all energy eigenstates rather than all eigenstates. Such a property is trivially true for local observables in translation invariant systems in the limit of infinite volume [BiroliWeakETH, MoriWeakETH, NormalWeakETH], for a translation invariant energy eigenbasis. Barring some trivial cases (such as noninteracting particles) where a highly degenerate spectrum ensures that (weak) eigenstate thermalization is not sufficient for thermalization in (almost) all physical states, this translation invariance property alone is sufficient for time-averaged thermalization in the absence of degeneracies, at least over infinitely long timescales.
@@
This is true even in “integrable” systems with an extensive number of local conserved quantities [BiroliWeakETH] (again for local observables in the infinite volume limit), and weak ETH has therefore been criticized (e.g., [deutsch2018eth]) as being of unclear relevance to thermalization. Here, we point out that this may be an effect of the infinite volume limit: the non-interacting ideal gas, despite being a prototypical “trivially” integrable system, is one of the few classical systems that can be rigorously proven to be ergodic and mixing (loosely, strongly thermalizing) in the infinite volume limit with fixed particle density [Sinai1976]. This can be understood as being due to the “diffusion” of information about the initial state to infinitely far away [GoldsteinLebowitzErgodicGas]. In such systems, it may be the case that more nonlocal observables that still remain sensitive to the volume of the system (such as the density of excitations, e.g., qubit “$1$” states, in a finite fraction of the total volume, coarse grained over a range of values to have only macroscopic accuracy) may be more nontrivial to consider, e.g., if one takes the thermodynamic limit by increasing the density of particles but keeping the volume fixed.
@@
A related objection to concluding thermalization from weak ETH is relevant in “quantum quench” protocols, where the initial state is prepared as a (low energy) stationary state of one Hamiltonian $\hat{H}_{0}$, while the dynamics implemented is that of a different Hamiltonian $\hat{H}$. Here, it is typically the case that the low energy states of $\hat{H}_{0}$ form (part of) the small fraction of states in which an observable satisfying weak ETH may fail to thermalize under the dynamics of $\hat{H}$ if the latter is “integrable” [RigolQuench]. For these quench protocols, we must develop a way to restrict the dynamics to an energy shell of $\hat{H}$ which has a large overlap with the low energy states of $\hat{H}_{0}$, though it is not presently clear how this can be achieved while maintaining rigorous accessibility [^22] As $\hat{H}_{0}$ does not in general commute with $\hat{H}$ by some significant amount, we do not expect that the echo strategy for conserved charges in Eq. (124) can be applied in a rigorous manner for similar reasons as in Eq. (19).. We also note that there are systems where the quench protocol leads to thermalization even with just weak ETH at higher energies [MoriThermalizationWithoutETH].
@@
Outside the context of quench protocols, however, such as in many-qubit systems where one expects to have the ability to prepare all computational basis states (for example), we expect that weak thermalization may be the strongest rigorous and accessible statement one can make in generic cases. This is indeed the case in classical statistical mechanics as well: one can almost never rule out a “measure” zero set of points such as periodic orbits (which are quite generic in “chaotic” systems [Haake, HOdA]) from failing to thermalize [Sinai1976, SinaiCornfeld, KhinchinStatMech]. Our view is that in systems where this property trivially follows for a certain class of observables due to some symmetry but does not describe the behavior of initial states of interest, one should either focus on a different class of observables, or restrict the Hilbert space to a smaller subspace with the initial states of interest, corresponding to the two cases described above.
@@ Acknowledgments
We thank Andrew Lucas for comments on the manuscript, Marcos Rigol for useful discussions on the physical relevance of “weak” eigenstate thermalization, and R. Shankar for providing a historical perspective on Ref. [JensenShankarETH]. This work was supported by the Heising-Simons Foundation under Grant 2024-4848.
@@ Appendix A Quantum Mazur-Suzuki inequality over finite times
$\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}\mathrm{d}t\ \langle A(t)A(0)\rangle\geq\sum_{k}\frac{\langle AQ_{k}\rangle^{2}}{\langle Q_{k}^{2}\rangle}.$ (174)
In finite-dimensional quantum systems, this has a significant issue with accessibility [ProsenMazur, DharMazur]: as with Eq. (12), this requires the infinite time average [SuzukiPhysica] to be over longer timescales than the energy level spacings, where each $Q_{k}=\sum_{n}q_{kn}\lvert E_{n}\rangle\langle E_{n}\rvert$ is a linear combination of a different subset of energy projectors (not necessarily spanning the full spectrum).
Here, we will show how completely positive averages of autocorrelators can be used to obtain a general analogue of Eq. (174) over finite timescales and even with approximate conserved quantities $\hat{Q}_{k}$ for any finite-dimensional quantum system. To our knowledge, this type of inequality has not previously been obtained from first principles in this general setting, but only in the classical limit [MazurPhysica] or in a strict thermodynamic limit for local systems, in the (weaker) limit of infinite times [ProsenMazur]. Starting with a time average [^23] For finite temperature averages with some density operator $\hat{\rho}(\hat{H})$, we note that all our considerations generalize if the inner product $\langle\hat{A},\hat{B}\rangle=\operatorname{Tr}[\hat{A}^{\dagger}\hat{B}]$ is replaced by $\langle\hat{A},\hat{B}\rangle_{\rho}=\operatorname{Tr}[\hat{\rho}(\hat{H})\hat{A}^{\dagger}\hat{B}]$ throughout, including in Eq. (175), provided that $\hat{\rho}(\hat{H})$ is exactly diagonal in the energy eigenbasis as per Eq. (19). In the context of the Mazur-Suzuki inequality, as our interest is not in rigorously determining the dynamics of other states in specific energy shells of interest as in Sec. 6, but only in the single state $\hat{\rho}(\hat{H})$ supported on the full system, it may be appropriate to regard the experimental inability to precisely prepare $\hat{\rho}(\hat{H})$ as a mere experimental error that changes the values of the measured correlators in the final inequality by some small $\epsilon$ compared to their ideal values. of the autocorrelator of $\hat{A}$ weighted by $w_{+}(t)$, and some energy band $\Delta E$ of interest where $\widetilde{w}_{+}(\delta E<\Delta E)>W$ (with $W<1$), we have from the derivation of Proposition 5.1 (see App. D.4.1):
$\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{A}(t)\hat{A}(0)]>W\operatorname{Tr}\left[[\hat{A}]_{\Delta E}^{2}\right].$ (175)
We recall that $[\hat{A}]_{\Delta E}$ has been defined in Eq. (7); by definition [^24] Here, we are making a subtle but unimportant switch to Hamiltonian evolution in the Heisenberg picture, in place of the Schrödinger picture in the rest of the text., $\hat{A}(t)=e^{i\hat{H}t}\hat{A}(0)e^{-i\hat{H}t}$ and $\hat{A}\equiv\hat{A}(0)$.
@@
$\geq\sum_{k=1}^{M_{Q}}\frac{\operatorname{Tr}[\hat{O}^{\dagger}\hat{Q}_{k}]\operatorname{Tr}[\hat{Q}_{k}^{\dagger}\hat{O}]}{\operatorname{Tr}[\hat{Q}_{k}^{\dagger}\hat{Q}_{k}]},$ (176)
@@
$\operatorname{Tr}\left([\hat{A}]_{\Delta E}^{2}\right)\geq\sum_{k=1}^{M_{Q}}\frac{\left(\operatorname{Tr}\left[[\hat{A}]_{\Delta E}\hat{Q}_{k}\right]\right)^{2}}{\operatorname{Tr}[\hat{Q}_{k}^{2}]}.$ (177)
Further noting that $\operatorname{Tr}\left([\hat{A}]_{\Delta E}\hat{Q}_{k}\right)=\operatorname{Tr}\left[\hat{A}[\hat{Q}_{k}]_{\Delta E}\right]$, this gives for the autocorrelator [with Eq. (175)]:
$\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{A}(t)\hat{A}(0)]>W\sum_{k=1}^{M_{Q}}\frac{\left(\operatorname{Tr}\left[\hat{A}[\hat{Q}_{k}]_{\Delta E}\right]\right)^{2}}{\operatorname{Tr}[\hat{Q}_{k}^{2}]}.$ (178)
@@
$\left\lvert\operatorname{Tr}[\hat{Q}_{k}^{\dagger}\hat{Q}_{j}]\right\rvert<\epsilon\sqrt{\operatorname{Tr}[\hat{Q}_{k}^{\dagger}\hat{Q}_{k}]\operatorname{Tr}[\hat{Q}_{j}^{\dagger}\hat{Q}_{j}]},$ (179)
for some small $\epsilon>0$. However, it is easy enough to account for this in practice without any nontrivial physics: we can derive a result for exactly orthogonal operators, and then consider the $O(\epsilon)$ errors made if the exactly orthogonal $\hat{Q}_{k}$ are replaced by experimentally accessible operators that are sufficiently close to them and satisfy Eq. (179) instead; we will therefore continue to assume exact orthogonality.
@@
$\delta Q_{k}^{2}[w_{2+}]\equiv\left\lvert\operatorname{Tr}[\hat{Q}_{k}^{2}]-\int\mathrm{d}t\ w_{2+}(t)\operatorname{Tr}[\hat{Q}_{k}(t)\hat{Q}_{k}(0)]\right\rvert.$ (180)
@@
$\implies\operatorname{Tr}[\hat{Q}_{k}^{2}]-\operatorname{Tr}\left([\hat{Q}_{k}]_{\Delta E}^{2}\right)=\operatorname{Tr}\left[\left(\hat{Q}_{k}-[\hat{Q}_{k}]_{\Delta E}\right)^{2}\right]$ $<\frac{\delta Q_{k}^{2}[w_{2+}]}{1-w_{20}}.$ (181)
@@
It follows that we can effectively replace $[\hat{Q}]_{\Delta E}$ in Eq. (178) with just $\hat{Q}_{k}$, making some small error. Specifically, we have by the Cauchy-Schwarz inequality,
@@
$\left\lvert\operatorname{Tr}\left[\hat{A}\hat{Q}_{k}\right]-\operatorname{Tr}\left[\hat{A}\ [\hat{Q}_{k}]_{\Delta E}\right]\right\rvert\leq\sqrt{\operatorname{Tr}[\hat{A}^{2}]\operatorname{Tr}\left[\left(\hat{Q}_{k}-[\hat{Q}_{k}]_{\Delta E}\right)^{2}\right]}<\sqrt{\frac{\delta Q_{k}^{2}[w_{2+}]}{1-w_{20}}\ \operatorname{Tr}[\hat{A}^{2}]}.$ (182)
@@
Using this in Eq. (178) with the triangle inequality [in the present context, the negative form $|x-y|\geq\left\lvert(|x|-|y|)\right\rvert$], which implies that
@@
$\left\lvert\operatorname{Tr}\left[\hat{A}\ [\hat{Q}_{k}]_{\Delta E}\right]\right\rvert\geq\left\lvert\left\lvert\operatorname{Tr}\left[\hat{A}\hat{Q}_{k}\right]\right\rvert-\sqrt{\frac{\delta Q_{k}^{2}[w_{2+}]}{1-w_{20}}\ \operatorname{Tr}[\hat{A}^{2}]}\right\rvert,$ (183)
@@
we obtain a rigorous finite-time Mazur-Suzuki inequality that applies to any finite-dimensional quantum system and accounts for approximate conserved quantities $\hat{Q}_{k}$ via Eq. (180):
@@
$\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{A}(t)\hat{A}(0)]>W\sum_{k}\frac{1}{\operatorname{Tr}[\hat{Q}_{k}^{2}]}\left\lvert\left\lvert\operatorname{Tr}\left[\hat{A}\hat{Q}_{k}\right]\right\rvert-\sqrt{\frac{\delta Q_{k}^{2}[w_{2+}]}{1-w_{20}}\ \operatorname{Tr}[\hat{A}^{2}]}\right\rvert^{2}.$ (184)
@@
This recovers Eq. (174) for infinite times and exact conserved quantities as $W\to 1$, $\delta Q_{k}^{2}[w_{2+}]\to 0$ and $w_{20}\to 0$. But we emphasize that in this formulation, despite being for a general finite-dimensional quantum system, it is entirely possible to take the thermodynamic limit $N\to\infty$ first before the infinite time limit [ProsenMazur].

### v2

@@ 8 Discussion: A sketch of experimental measurement protocols
If the $t_{j}$ are sampled irregularly with average spacing $\delta t$ according to some sufficiently random statistical (e.g. Poisson) process, then $\widetilde{v}(E)$ instead becomes quasiperiodic, with a significantly larger “recurrence” energy $E_{\text{rec}}$ (this is connected to quantum recurrence times [79, 80] for an irregular energy spectrum, except in our case time and energy have switched places). In particular, if one has $\tau$ such irregular samples of time and $\widetilde{v}(E_{0})=1$, one expects $\widetilde{v}(E)\ll 1$ between $E_{c}$ and $E_{c}\pm E_{\text{rec}}$, where
@@
This means that the number of “extraneous” energy bands or shells is reduced exponentially with the number of samples $\tau$ if one adopts an irregular sampling of times. Often, the “full width” of the energy spectrum in a many-body system (e.g., the separation between the lowest and highest energy levels) scales [81, 82] as a small power of $N$:
@@
A remaining question is whether we can still generate completely positive functions $v_{+}(t)$ of this type. This is indeed possible [^20] For a related argument in a separate context that involves convolutions and applies to real-valued functions, due to Laura Shou, see Ref. [28, Appendix D 1]. We use a slightly different argument here to avoid the requirement of real-valuedness, i.e., we do not require $\mathbbR\sum_{k=1}^{\eta}e^{-i\zeta_{k}(E-E_{c})}\in\mathbb{R}$, which allows a more unconstrained choice of the $\zeta_{k}$., and one example is the spectral form factor $K(t)=\lvert L_{\mathcal{H}}(t)\rvert^{2}$ of some energy spectrum, which remains non-negative $K(t)>0$, and whose Fourier transform is the probability distribution (and therefore, non-negative) of two-level spacings in the spectrum. In our case, where the samples are of time rather than energy levels, we should identify an analogue of the spectral form factor with the energy domain and the $2$-level distribution with the time domain. This suggests that, given a set of $\eta$ points $\zeta_{1},\ldots,\zeta_{\eta}$, we can set $\tau=\eta^{2}$ and
@@ 9 Conclusion
We have described an approach to quantum statistical mechanics that is based on the accessible dynamical properties of a single initial state corresponding to an observable instead of the detailed properties of the energy levels, mirroring classical approaches [8, 9] that bypass the ergodic hypothesis (see also [83] for a historical survey of related issues). Crucially, as described in Sec. 8, this allows us to make a conclusive experimental determination of thermalization even in large systems of several qubits with tractable resources. While our results center around energy-band thermalization as a finite-resolution variant of eigenstate thermalization, which is aesthetically desirable given the viewpoint that the energy eigenvalues and eigenstates completely determine quantum dynamics, we found that it is possible, and perhaps even more convenient, to directly connect time-domain quantities to thermalization and bypass the energy domain [^21] We view this as being analogous to the “energy-time uncertainty principle” for many-body systems in Refs. [27, 28], where instead of using the spectral form factor to determine the structure of the energy spectrum and then formulate a speed limit in terms of energy parameters, we directly formulate a time-domain speed limit in terms of the time-domain spectral form factor, which implicitly contains information about the energy domain but is best measured in the time domain..
From an analytical standpoint, our results generalize the observed connection between classical ergodicity and eigenstate thermalization in quantized classical systems [42, 43, 44, 40, 41, 45, 46, 47] to observable-dependent quantum thermalization in fully quantum systems. In doing so, we have introduced a framework to account for thermalization over finite energy and time scales, in particular showing that these finite-time features are sufficient to determine thermalization over arbitrarily long timescales due to interference effects (which, to our knowledge, has no obvious classical counterpart). In addition to thermalization, we used this approach to obtain a finite-time Mazur-Suzuki inequality for autocorrelators in terms of approximately conserved charges in App. A, which appears to be the first rigorous result of this nature and may be useful for quantum transport problems in finite but large systems. We have also developed intrinsically quantum methods to access thermalization in energy shells and conserved charges that again rely on interference effects with no classical counterpart.
@@
Finally, we address an important question that has been of some concern in the literature, which is the problem of the physical relevance of “weak” thermalization. In particular, our results show that thermalization can be established in an accessible manner for “almost all” (rather than precisely all) physical states (i.e., “weak” thermalization), such as computational basis states. While our methods can also show thermalization in every conceivable initial state (i.e., “strong” thermalization), as discussed in Sec. 4, this requires a level of accuracy that we do not expect to be experimentally accessible (nor analytically tractable except perhaps in special cases [51]) in the thermodynamic limit of many particles. Given that our focus has been on “accessible” aspects of quantum statistical mechanics, it is worth highlighting some common objections to weak thermalization and their relevance to our approach.
@@
Let us then consider “weak ETH”, which refers to the diagonal part of ETH [Eq. (3)] applying to almost all energy eigenstates rather than all eigenstates. Such a property is trivially true for local observables in translation invariant systems in the limit of infinite volume [34, 35, 84], for a translation invariant energy eigenbasis. Barring some trivial cases (such as noninteracting particles) where a highly degenerate spectrum ensures that (weak) eigenstate thermalization is not sufficient for thermalization in (almost) all physical states, this translation invariance property alone is sufficient for time-averaged thermalization in the absence of degeneracies, at least over infinitely long timescales.
@@
This is true even in “integrable” systems with an extensive number of local conserved quantities [34] (again for local observables in the infinite volume limit), and weak ETH has therefore been criticized (e.g., [19]) as being of unclear relevance to thermalization. Here, we point out that this may be an effect of the infinite volume limit: the non-interacting ideal gas, despite being a prototypical “trivially” integrable system, is one of the few classical systems that can be rigorously proven to be ergodic and mixing (loosely, thermalizing) in the infinite volume limit with fixed particle density [7]. This can be understood as being due to the “diffusion” of information about the initial state to infinitely far away [85, 86]. In such systems, it may be the case that more nonlocal observables that still remain sensitive to the volume of the system (such as a current density of excitations e.g. average “hopping” velocity of qubit “11” states in a finite fraction of the total volume, coarse grained over a range of values to have only macroscopic accuracy) may be more nontrivial to consider, e.g., if one takes the thermodynamic limit by increasing the density of particles but keeping the volume fixed.
@@
A related objection to concluding thermalization from weak ETH is relevant in “quantum quench” protocols, where the initial state is prepared as a (low energy) stationary state of one Hamiltonian $\hat{H}_{0}$, while the dynamics implemented is that of a different Hamiltonian $\hat{H}$. Here, it is typically the case that the low energy states of $\hat{H}_{0}$ form (part of) the small fraction of states in which an observable satisfying weak ETH may fail to thermalize under the dynamics of $\hat{H}$ if the latter is “integrable” [87]. For these quench protocols, we must develop a way to restrict the dynamics to an energy shell of $\hat{H}$ which has a large overlap with the low energy states of $\hat{H}_{0}$, though it is not presently clear how this can be achieved while maintaining rigorous accessibility [^22] As $\hat{H}_{0}$ does not in general commute with $\hat{H}$ by some significant amount, we do not expect that the echo strategy for conserved charges in Eq. (124) can be applied in a rigorous manner for similar reasons as in Eq. (19).. We also note that there are systems where the quench protocol leads to thermalization even with just weak ETH at higher energies [88].
@@
Outside the context of quench protocols, however, such as in many-qubit systems where one expects to have the ability to prepare all computational basis states (for example), we expect that weak thermalization may be the strongest rigorous and accessible statement one can make in generic cases (e.g., an analysis in terms of Turing machines suggests that for translation-invariant Hamiltonians, thermalization in specific initial states is a computationally undecidable problem [89, 90], translating in our case to $\epsilon$ in Theorem 4.3 being insufficiently small for thermalization in all states). This is indeed the case in classical statistical mechanics as well: one can almost never rule out a “measure” zero set of points such as periodic orbits (which are quite generic in “chaotic” systems [62, 91]) from failing to thermalize [7, 6, 8]. Our view is that in systems where this property trivially follows for a certain class of observables due to some symmetry but does not describe the behavior of initial states of interest, one should either focus on a different class of observables, or restrict the Hilbert space to a smaller subspace with the initial states of interest, corresponding to the two cases described above.
@@ Acknowledgments
We thank Andrew Lucas for comments on the manuscript, Marcos Rigol for useful discussions on the physical relevance of “weak” eigenstate thermalization, and R. Shankar for providing a historical perspective on Ref. [11]. This work was supported by the Heising-Simons Foundation under Grant 2024-4848.
@@ Appendix A Quantum Mazur-Suzuki inequality over finite times
The Mazur-Suzuki inequality [48, 49] states that for an observable $A$ in a classical or quantum system with exact “orthogonal” conserved quantities $Q_{k}$ and expectation values $\langle\cdot\rangle$ (so that $\langle Q_{k}Q_{j}\rangle=\langle Q_{k}^{2}\rangle\delta_{kj}$), the infinite time average of the autocorrelator is bounded by:
$\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}\mathrm{d}t\ \langle A(t)A(0)\rangle\geq\sum_{k}\frac{\langle AQ_{k}\rangle^{2}}{\langle Q_{k}^{2}\rangle}.$ (A.1)
In finite-dimensional quantum systems, this has a significant issue with accessibility [53, 54]: as with Eq. (12), this requires the infinite time average [49] to be over longer timescales than the energy level spacings, where each $Q_{k}=\sum_{n}q_{kn}\lvert E_{n}\rangle\langle E_{n}\rvert$ is a linear combination of a different subset of energy projectors (not necessarily spanning the full spectrum).
@@
$\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{A}(t)\hat{A}(0)]>W\operatorname{Tr}\left[[\hat{A}]_{\Delta E}^{2}\right].$ (A.2)
We recall that $\hat{A}]_{\Delta E}$ has been defined in Eq. (7); by definition [^24] Here, we are making a subtle but unimportant switch to Hamiltonian evolution in the Heisenberg picture, in place of the Schrödinger picture in the rest of the text., $\hat{A}(t)=e^{i\hat{H}t}\hat{A}(0)e^{-i\hat{H}t}$ and $\hat{A}\equiv\hat{A}(0)$.
@@
$\geq\sum_{k=1}^{M_{Q}}\frac{\operatorname{Tr}[\hat{O}^{\dagger}\hat{Q}_{k}]\operatorname{Tr}[\hat{Q}_{k}^{\dagger}\hat{O}]}{\operatorname{Tr}[\hat{Q}_{k}^{\dagger}\hat{Q}_{k}]},$ (A.3)
@@
$\operatorname{Tr}\left([\hat{A}]_{\Delta E}^{2}\right)\geq\sum_{k=1}^{M_{Q}}\frac{\left(\operatorname{Tr}\left[[\hat{A}]_{\Delta E}\hat{Q}_{k}\right]\right)^{2}}{\operatorname{Tr}[\hat{Q}_{k}^{2}]}.$ (A.4)
Further noting that $\operatorname{Tr}\left([\hat{A}]_{\Delta E}\hat{Q}_{k}\right)=\operatorname{Tr}\left[\hat{A}[\hat{Q}_{k}]_{\Delta E}\right]$, this gives for the autocorrelator [with Eq. (A.2)]:
$\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{A}(t)\hat{A}(0)]>W\sum_{k=1}^{M_{Q}}\frac{\left(\operatorname{Tr}\left[\hat{A}[\hat{Q}_{k}]_{\Delta E}\right]\right)^{2}}{\operatorname{Tr}[\hat{Q}_{k}^{2}]}.$ (A.5)
@@
$\left\lvert\operatorname{Tr}[\hat{Q}_{k}^{\dagger}\hat{Q}_{j}]\right\rvert<\epsilon\sqrt{\operatorname{Tr}[\hat{Q}_{k}^{\dagger}\hat{Q}_{k}]\operatorname{Tr}[\hat{Q}_{j}^{\dagger}\hat{Q}_{j}]},$ (A.6)
for some small $\epsilon>0$. However, it is easy enough to account for this in practice without any nontrivial physics: we can derive a result for exactly orthogonal operators, and then consider the $O(\epsilon)$ errors made if the exactly orthogonal $\hat{Q}_{k}$ are replaced by experimentally accessible operators that are sufficiently close to them and satisfy Eq. (A.6) instead; we will therefore continue to assume exact orthogonality.
@@
$\delta Q_{k}^{2}[w_{2+}]\equiv\left\lvert\operatorname{Tr}[\hat{Q}_{k}^{2}]-\int\mathrm{d}t\ w_{2+}(t)\operatorname{Tr}[\hat{Q}_{k}(t)\hat{Q}_{k}(0)]\right\rvert$. (A.7)
@@
$\implies\operatorname{Tr}[\hat{Q}_{k}^{2}]-\operatorname{Tr}\left([\hat{Q}_{k}]_{\Delta E}^{2}\right)=\operatorname{Tr}\left[\left(\hat{Q}_{k}-[\hat{Q}_{k}]_{\Delta E}\right)^{2}\right]$ $<\frac{\delta Q_{k}^{2}[w_{2+}]}{1-w_{20}}.$ (A.8)
@@
It follows that we can effectively replace $\hat{Q}]_{\Delta E}$ in Eq. (A.5) with just $\hat{Q}_{k}$, making some small error. Specifically, we have by the Cauchy-Schwarz inequality,
@@
$\left\lvert\operatorname{Tr}\left[\hat{A}\hat{Q}_{k}\right]-\operatorname{Tr}\left[\hat{A}\ [\hat{Q}_{k}]_{\Delta E}\right]\right\rvert\leq\sqrt{\operatorname{Tr}[\hat{A}^{2}]\operatorname{Tr}\left[\left(\hat{Q}_{k}-[\hat{Q}_{k}]_{\Delta E}\right)^{2}\right]}<\sqrt{\frac{\delta Q_{k}^{2}[w_{2+}]}{1-w_{20}}\ \operatorname{Tr}[\hat{A}^{2}]}.$ (A.9)
@@
Using this in Eq. (A.5) with the triangle inequality [in the present context, the negative form $|x-y|\geq\left\lvert(|x|-|y|)\right\rvert$], which implies that
@@
$\left\lvert\operatorname{Tr}\left[\hat{A}\ [\hat{Q}_{k}]_{\Delta E}\right]\right\rvert\geq\left\lvert\left\lvert\operatorname{Tr}\left[\hat{A}\hat{Q}_{k}\right]\right\rvert-\sqrt{\frac{\delta Q_{k}^{2}[w_{2+}]}{1-w_{20}}\ \operatorname{Tr}[\hat{A}^{2}]}\right\rvert$, (A.10)
@@
we obtain a rigorous finite-time Mazur-Suzuki inequality that applies to any finite-dimensional quantum system and accounts for approximate conserved quantities $\hat{Q}_{k}$ via Eq. (A.7):
@@
$\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{A}(t)\hat{A}(0)]>W\sum_{k}\frac{1}{\operatorname{Tr}[\hat{Q}_{k}^{2}]}\left\lvert\left\lvert\operatorname{Tr}\left[\hat{A}\hat{Q}_{k}\right]\right\rvert-\sqrt{\frac{\delta Q_{k}^{2}[w_{2+}]}{1-w_{20}}\ \operatorname{Tr}[\hat{A}^{2}]}\right\rvert^{2}.$ (A.11)
@@
This recovers Eq. (A.1) for infinite times and exact conserved quantities as $W\to 1$, $\delta Q_{k}^{2}[w_{2+}]\to 0$ and $w_{20}\to 0$. But we emphasize that in this formulation, despite being for a general finite-dimensional quantum system, it is entirely possible to take the thermodynamic limit $N\to\infty$ first before the infinite time limit [53].